What is the volume in (mL) of 2.17 x 1024 molecules dihydrogen monoxide?
1
Expert's answer
2020-09-15T13:49:38-0400

2.17 x 1024molecules dihydrogen monoxide .


MOLES = 2.17 * 10^24 / 6.02 * 10^23 .

= 2.17 / 6.023 * 10

= 3.60 MOLE .


Volume =( 3.60 * 22.4 ) liter = 80.64 liter = ( 80.64 * 1000 ) ml . = 80640 ml .